vidyo

vidyo link

Action Items:

News:

* from CMS week:

  • upgrade ->
    • work on TDRs on-going. First tracker, then muons.
    • Sandro is working to develop solution to reduce cost of DT upgrade.
    • riunione commissione 1 today and tomorrow -> discussion on upgrade and on new timing device
  • CB ->
    • main objections to new timing device came from camporesi and sphicas (new high gran ECAL)
    • guidelines on how to write a paper have been produced
    • raised question -> one week needed to pass from green light to paper submission (atlas takes 1 day only). additional man power will be set.
    • 171 nominations for new convenerships starting from Sept (roberto, franco, martino m, marco, paul nominated)
    • EPR - 199 institutions, about 100 have done more than their duty. Padova did 91% (20 inst with >90%). Most of our pledges from legnaro (sistemisti). Other contributions from Anna, Jacopo and Lisa, Alessio. They will probably not be able to replicate this year. Please subscribe to shifts and check for contacts position!

* post-doc INFN per stranieri:

  • not officially, our best 'candidate' is Tico (ttH full hadronic) on 23rd position. 20 positions open.

* summer student project -> patrizia is not available this summer. Paul will be. Project to be submitted within 1 week.

Framework e produzione

  • new imags created today using a new file system which is not lustre. you do not have your home but you can run an elastic cluster...
  • old datasets have been deleted in legnaro. please continue to keep it clean in the future. 23TB of free space now.
  • all re-reco + almost all Summer16 are in Legnaro.
  • alpha framework have been updated to moriond17 and re-reco. production for diboson is currently on-going. lsf resources are now packed. HH and same-sign production will come later.
  • paolo is going to produce ntuples for b-physics. probably help needed.
  • new Thursday presentation from Marco about big data (infrastucture) project proposed to EU

Attivita' di analisi

B physics

  • Stefano will not accept convenership nomination, probably next
  • P5' presentation last week at physics week. next presentation at PAG tomorrow. Fit finished with some problems in 2 bins (high chi2). we hope to go for moriond.
  • nicola and roberto are developing new HLT paths based on one L1 only. first plots this morning, deadline this Thurs.

Dark Matter

  • SewYan has resumed work on that. Not targeting moriond. he should go to fermilab within this month.
  • student from astronomy asked a thesis on DM to Marco. End of this year.

Dibosoni, VZ risonante (V=W,Z)

  • Jacopo will not have time anymore to work 100% on the analysis.
  • Lisa took over the ntuple production -> ntuple unified between the two channels in order to have a single production step.
  • invisible -> B2G-17-005 (only invisible)
    • B2G conveners asked for object review within this week. Pre-appr in 1 week.
    • production on-going.
    • out from moriond EW but on track for moriond qcd.
    • a paper on invisible only will be probably allowed.

hh4b non risonante

  • not anymore on moriond rush since result on ichep dataset was not satisfying (not showed even to pag)
  • from discussion with convener about hh4b paper strategy:
    • non-resonant will go with a standalone paper on 2016data as soon as a reasonable result is ready.
    • resonant resolved-boosted need instead to be probably combined even if they have 2 cadi lines in different PAGs (resolved has not been pre-approved last week).
  • our strategy is to wok on moriond datasets and to try to improve/change fit strategy.
  • pre-approval of other decay channels are on-going (bbtautau set an expected limit equal to 29xSM)

Same sign leptons

  • v4 of AN delivered 10 days ago with answers to raised points
  • results form other analysis do not confirm mine. They are probably over estimating the systematic. Paolo-> we need to rise hands and officially tell something about it.
  • Ugo will circulate this PAS to everybody. Please check it.

Seesaw a 13 TeV

  • PAS has been presented.

EPR

Isolamento muoni

  • Lisa -> POG asked to update iso for moriond but it was impossible due to schedule.

Flavor tagging

  • Alessio is not working anymore on charge tagger. Documentation is needed and task will be done then

Trigger

Detector

A.O.B.:

-- MarcoZanetti - 2016-02-01

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2017-02-06 - MartinoDallOsso
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback